# Zero-sum partitions of Abelian groups and their applications to magic- and antimagic-type labelings

### v4

@@ 3. Some Applications
3.1. $\Gamma$-irregular labeling for digraphs
@@
Lemma 3.1 ([8]).
A digraph $\overrightarrow{G}=(V,A)$ with no isolated vertices has a $\Gamma$-irregular labeling if and only if there exists an injective mapping $\varphi$ from $V$ to $\Gamma$ such that $\sum_{x\in V(\overrightarrow{C})}\varphi(x)=0$ for every weakly connected component $\overrightarrow{C}$ of $\overrightarrow{G}$.
@@
3.2. $\Gamma$-antimagic labeling
Let $\Gamma$ be an Abelian group. A $\Gamma$-antimagic labeling of a graph $G$ is defined as a bijection $f\colon E(G)\to\Gamma$, where the weight of each vertex, which is the sum of the labels on its incident edges, is distinct for all vertices. Kaplan et al. [16] demonstrated that if $\Gamma$ has a unique involution, then no tree of order $|\Gamma|$ has a $\Gamma$-antimagic labeling. They conjectured that a tree of order $|\Gamma|$ has a $\Gamma$-antimagic labeling if and only if $\Gamma$ does not have a unique involution. By applying the same approach used in [16] for 2-trees, and using Corollary 2.6, we derive the following result:
@@ 3.2. $\Gamma$-antimagic labeling
The necessity of the condition follows directly from the work of Kaplan et al. (see Theorem E, [16]). Thus, let us assume that $|I(\Gamma)|\neq 1$. Let $\{v_{i}\}_{i=1}^{t}$ be the vertices of $T$ which are not leaves. Let us denote their corresponding numbers of children by $\{m_{i}\}_{i=1}^{t}$. Thus $\sum_{i=1}^{t}m_{i}=|\Gamma|-1$. Since $T$ is a $4$-tree, we have that $m_{i}\geq 4$ for every $1\leq i\leq t$. The group $\Gamma$ has $4$-ZSPP by Corollary 2.6, and there exists a zero-sum partition $\{S_{i}\}_{i=1}^{t}$ of $\Gamma^{*}$. For every $1\leq i\leq t$, we label the edges of the set $E_{i}=\{v_{i}w\colon w$ is a child of $v_{i}\}$ by the elements of $S_{i}$. The edges of $T$ are labeled bijectively with the non-zero elements of $\Gamma$, and the sum of the labels in every $E_{i}$ is 00. Since every vertex of $T$, except the root (whose weight is 00), has a unique parent, it easily follows that the weights are pairwise distinct. ∎
3.3. $\Gamma$-distance-magic and -antimagic labeling in graphs with twins
Froncek [12] defined the notion of a group-distance-magic labeling of a graph $G=(V,E)$. In such a labeling, the vertices of the graph are labeled through a bijection with the elements of an Abelian group $\Gamma$. The weight of each vertex is computed as the sum (in $\Gamma$) of the labels assigned to its neighbors. If all weights are the same, then it is a $\Gamma$-distance-magic labeling. If all weights are different, then it is a $\Gamma$-distance-antimagic labeling [1].
@@ 3.3. $\Gamma$-distance-magic and -antimagic labeling in graphs with twins
Notice that Corollary 3.10 extends the results on $\Gamma$-distance-magic labelings of complete $k$-partite graphs presented in [7].
@@
For the last two results of this subsection, let us recall the definition of a $G$-join introduced by Sabidussi [19].
Definition 3.12 ([19]).
@@
Note that the equivalence classes of twins in a graph can be found in linear time [14]. So, if a respective zero-sum partition of $\Gamma^{*}$ is given or it can be found efficiently, then we can efficiently compute the corresponding labeling.
@@ 4. Final Remarks
Finally, let us complement Definition 1.4. Let $S$ be a subset of cardinality $m=6k+s$, for a positive integer $k$ and $s\in\{0,2,4\}$, of a finite Abelian group $\Gamma$. A partition of $S$ into $k$ good $6$-subsets and $s/2$ zero-sum $2$-subsets is called a Skolem partition of $S$. It is known that $\mathbb{Z}_{m}^{*}\setminus I(\mathbb{Z}_{m})$, for $m=6k+s$ with any positive integer $k$, has a Skolem partition when $s\in\{0,4\}$, but the situation is slightly different for $s=2$. For $m\equiv 2$ or $8\pmod{24}$, there exists a Skolem partition, whereas for $m\equiv 14$ or $20\pmod{24}$, such a partition does not exist (see Lemma 4 in [22]). Note that there exist groups $\Gamma$ with $|I(\Gamma)|>1$ such that the set $R=\Gamma^{*}\setminus I(\Gamma)$ has a Skolem partition. For instance, let $\Gamma\cong(\mathbb{Z}_{2})^{\eta}\times H$ for some natural number $\eta>1$ and $|H|\equiv 1\pmod{6}$. By Theorem 1.5, there exists a partition of $H^{*}$ into good $6$-subsets:
@@
$R=\left(\bigcup\limits_{a\in L}\bigcup\limits_{i=1}^{|H^{*}|/6}\{(a,b_{i}),(\phi(a),c_{i}),(\varphi(a),-b_{i}-c_{i}),(-a,-b_{i}),(-\phi(a),-c_{i}),(-\varphi(a),b_{i}+c_{i})\}\right).$
@@
As reflected in Proposition 3.8, given a graph $G=(V,E)$ with a partition of the vertex set into subsets of false twins, existence of a zero-sum partition of an Abelian group $\Gamma$ into subsets of corresponding cardinalities is a sufficient condition for existence of a $\Gamma^{*}$-distance-magic labeling of $G$. One can easily see that this condition is not necessary. For example, consider the group $\mathbb{Z}_{6}$. There exists no zero-sum partition of $\mathbb{Z}_{6}^{*}$ as required in Proposition 3.8. On the other hand, $\{\{1,2\},\{3\},\{4,5\}\}$ is a constant-sum partition of $\mathbb{Z}^{*}_{6}$, with constant sum $3$, which leads to a $\mathbb{Z}_{6}^{*}$-distance-magic labeling of the complete tripartite graph $K_{1,2,2}$. So far, constant-sum partitions of Abelian groups have been studied only for the case of partitioning into three subsets (see [5]). It would be interesting to consider constant-sum partitions into more than three subsets.
@@
5. Statements and Declarations
  • Funding. The work of the first author was supported by the program “Excellence initiative – research university” for the AGH University.

### v5

@@ 3 Some Applications
3.1 $\Gamma$-irregular labeling for digraphs
@@
Lemma 3.1 ((Cichacz and Tuza, 2022)).
A digraph $\overrightarrow{G}=(V,A)$ with no isolated vertices has a $\Gamma$-irregular labeling if and only if there exists an injection $\varphi$ from $V$ to $\Gamma$ such that $\sum_{x\in V(\overrightarrow{C})}\varphi(x)=0$ for every weakly connected component $\overrightarrow{C}$ of $\overrightarrow{G}$.
@@
3.2 $\Gamma$-antimagic labeling
Let $\Gamma$ be an Abelian group. A $\Gamma$-antimagic labeling of a graph $G$ is defined as a bijection $f\colon E(G)\to\Gamma$, where the weight of each vertex, which is the sum of the labels on its incident edges, is distinct for all vertices. Kaplan et al. (2009) demonstrated that if $\Gamma$ has a unique involution, then no tree of order $|\Gamma|$ has a $\Gamma$-antimagic labeling. They conjectured that a tree of order $|\Gamma|$ has a $\Gamma$-antimagic labeling if and only if $\Gamma$ does not have a unique involution. By applying the same approach used by Kaplan et al. (2009) for 2-trees, and using Corollary 2.6, we derive the following result:
@@ 3.2 $\Gamma$-antimagic labeling
The necessity of the condition follows directly from the work of Kaplan et al. (2009, Theorem E). Thus, let us assume that $|I(\Gamma)|\neq 1$. Let $\{v_{i}\}_{i=1}^{t}$ be the vertices of $T$ which are not leaves. Let us denote their corresponding numbers of children by $\{m_{i}\}_{i=1}^{t}$. Thus $\sum_{i=1}^{t}m_{i}=|\Gamma|-1$. Since $T$ is a $4$-tree, we have that $m_{i}\geq 4$ for every $1\leq i\leq t$. The group $\Gamma$ has $4$-ZSPP by Corollary 2.6, and there exists a zero-sum partition $\{S_{i}\}_{i=1}^{t}$ of $\Gamma^{*}$ such that $|S_{i}|=m_{i}$ for $1\leq i\leq t$. For every $1\leq i\leq t$, we label the edges of the set $E_{i}=\{v_{i}w\colon w$ is a child of $v_{i}\}$ by the elements of $S_{i}$. The edges of $T$ are labeled bijectively with the non-zero elements of $\Gamma$, and the sum of the labels in every $E_{i}$ is 00. Since every vertex of $T$, except the root (whose weight is 00), has a unique parent, it easily follows that the weights are pairwise distinct. ∎
3.3 $\Gamma$-distance-magic and -antimagic labeling in graphs with twins
Froncek (2013b) defined the notion of a group-distance-magic labeling of a graph $G=(V,E)$. In such a labeling, the vertices of the graph are labeled through a bijection with the elements of an Abelian group $\Gamma$. The weight of each vertex is computed as the sum (in $\Gamma$) of the labels assigned to its neighbors. If all weights are the same, then it is a $\Gamma$-distance-magic labeling. If all weights are different, then it is a $\Gamma$-distance-antimagic labeling (Anholcer et al., 2021).
@@ 3.3 $\Gamma$-distance-magic and -antimagic labeling in graphs with twins
Notice that Corollary 3.10 extends the results on $\Gamma$-distance-magic labelings of complete $k$-partite graphs presented by Cichacz and Suchan (2023).
@@
For the last two results of this subsection, let us recall the definition of a $G$-join introduced by Sabidussi (1961).
Definition 3.12 ((Sabidussi, 1961)).
@@
Note that the equivalence classes of twins in a graph can be found in linear time (Habib and Paul, 2010). So, if a respective zero-sum partition of $\Gamma^{*}$ is given or it can be found efficiently, then we can efficiently compute the corresponding labeling.
@@ 4 Final Remarks
Finally, let us complement Definition 1.4. Let $S$ be a subset of cardinality $m=6k+s$, for a positive integer $k$ and $s\in\{0,2,4\}$, of a finite Abelian group $\Gamma$. A partition of $S$ into $k$ good $6$-subsets and $s/2$ zero-sum $2$-subsets is called a Skolem partition of $S$. It is known that $\mathbb{Z}_{m}^{*}\setminus I(\mathbb{Z}_{m})$, for $m=6k+s$ with any positive integer $k$, has a Skolem partition when $s\in\{0,4\}$, but the situation is slightly different for $s=2$. For $m\equiv 2$ or $8\pmod{24}$, there exists a Skolem partition, whereas for $m\equiv 14$ or $20\pmod{24}$, such a partition does not exist (Tannenbaum, 1981, Lemma 4)). Note that there exist groups $\Gamma$ with $|I(\Gamma)|>1$ such that the set $R=\Gamma^{*}\setminus I(\Gamma)$ has a Skolem partition. For instance, let $\Gamma\cong(\mathbb{Z}_{2})^{\eta}\times H$ for some natural number $\eta>1$ and $|H|\equiv 1\pmod{6}$. By Theorem 1.5, there exists a partition of $H^{*}$ into good $6$-subsets:
@@
$R=\bigcup\limits_{a\in L}\bigcup\limits_{i=1}^{|H^{*}|/6}\{(a,b_{i}),(\phi(a),c_{i}),(\varphi(a),-b_{i}-c_{i}),(-a,-b_{i}),(-\phi(a),-c_{i}),(-\varphi(a),b_{i}+c_{i})\}.$
@@
As reflected in Proposition 3.8, given a graph $G=(V,E)$ with a partition of the vertex set into subsets of false twins, existence of a zero-sum partition of an Abelian group $\Gamma$ into subsets of corresponding cardinalities is a sufficient condition for existence of a $\Gamma^{*}$-distance-magic labeling of $G$. One can easily see that this condition is not necessary. For example, consider the group $\mathbb{Z}_{6}$. There exists no zero-sum partition of $\mathbb{Z}_{6}^{*}$ as required in Proposition 3.8. On the other hand, $\{\{1,2\},\{3\},\{4,5\}\}$ is a constant-sum partition of $\mathbb{Z}^{*}_{6}$, with constant sum $3$, which leads to a $\mathbb{Z}_{6}^{*}$-distance-magic labeling of the complete tripartite graph $K_{1,2,2}$. So far, constant-sum partitions of Abelian groups have been studied only for the case of partitioning into three subsets (Cichacz, 2017). It would be interesting to consider constant-sum partitions into more than three subsets.
@@
5 Acknowledgments
The authors would like to thank the anonymous reviewers for their careful reading of the manuscript and their insightful comments and suggestions.
6 Statements and Declarations
  • Funding. This work was supported by the program “Excellence initiative – research university” for the AGH University.